five kiwi fruit and four apples cost $2.06. A kiwi fruit and five apples cost $1.21. how many cents does a kiwi fruit cost?
Found 2 solutions by josgarithmetic, jackiewalker123:
Answer by josgarithmetic(39792)

k = cents per kiwi
p = center per aPple

5 kiwis and 4 apples, 5k+4p=206
1 kiwi and 5 apples, k+5p=121

Two linear equations in two variables.

Answer by jackiewalker123(2)

k = cents per kiwi
p = center per aPple
5 kiwis and 4 apples, 5k+4p=206
1 kiwi and 5 apples, k+5p=121
5k+4p=206 -->(1)
k+5p=121-->(2)
Multiply (1) by 5 and (2) by 4 we get
25k+20p= 1030 (3)
4k+20p=484 (4)
Subtract =>21k =546
K= 26
sub.into (2) => 26+5p=121
5p= 95
P= 19
solution: kiwi fruit cost k=26 cents.